Given the reaction below, what is the mass of the products formed when 30 liters of CO gas is used at

STP?


I 2 O 5 (s) + CO(g) I 2 (s) + CO 2 (g)


1
Expert's answer
2021-02-18T07:46:16-0500

I2O5(s) + 5CO(g) → I2(s) + 5CO2(g)

1 mole of any gas occupies 22.4 L at STP.

Proportion:

1 mol – 22.4 L

x – 30 L

"x = \frac{1 \times 30}{22.4}=1.34 \;mol"

n(CO) = 1.34 mol

According to the reaction:

n(I2) "= \frac{n(CO)}{5} = \frac{1.34}{5} = 0.268 \;mol"

"m = n \times M"

M(I2) = 253.8 g/mol

m(I2) "= 0.268 \times 253.8 = 68.02 \;g"

According to the reaction:

n(CO2) = n(CO) = 1.34 mol

M(CO2) = 44.01 g/mol

m(CO2) "= 44.01 \times 1.34 = 58.97 \;g"